Substitute y = 4 into the following equation: 4x + 2y = 52. What is the value of x?

Answer by ewatrrr(24785)

4x + 2y = 52
Substitute y = 4
4x + 2(4) = 52
4x + 8 = 52
4x = 44
x = 11